A refrigerator has a COP of 5.5 how much work is needed to make ice cubes from 1 liter of water at 10 degrees celcius? hint: 1 liter of water = how many kilograms? c= 4184 jouls / kgkelvin
latent heat of fusion os water is 333500 j/kg
Answer #1

4 m QE # Heat to be extracted, QE = mc(OT) mlf 4184 (10-0) + 33350 Qe = ) soo 1 375340 # QE COP W - S.5 375340 И) Work, W = 6